OpenStudy (anonymous):

Integral of x^5*ln(3x)dx

OpenStudy (watchmath):

\(u=\ln(3x)\) and \(dv=x^5\,dx\) Then \(du=dx/x and v=x^6/6\) It follows your integral equals \(=x^6\ln(3x)/6-\int x^5/6\, dx\) \(=x^6\ln(3x)/6-x^6/36 +C\)

OpenStudy (anonymous):

What is this question in relation to?